The cashier's office at the college has collected $7020 from part-time students for a total of 42 courses. If each credit hour costs $45, how many 3-hour courses and how many 5-hour courses were paid for?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

T+F=42 OR T=42-F
45*3T+45*5F=7020 NOW SUBSTITUTE (42-F) FOR T & SOLVE FOR F
45*3(42-F)+45*5F=7020
135(42-F)+225F=7020
5670-135F+225F=7020
90F=7020-5670
90F=1350
F=1350/90
F=15 NUMBER OF 5 CREDIT COURSES.
T=42-15
T=27 NUMBER OF 3 CREDIT COURSES.
PROOF
45*3*27+45*5*15=7020
3645+3375=7020
7020=7020

Answer by stanbon(75887)

The cashier's office at the college has collected $7020 from part-time students for a total of 42 courses. If each credit hour costs $45, how many 3-hour courses and how many 5-hour courses were paid for?
------------
Let # of 3-hr. courses be "x"; income from these is 3*45*x = 135x dollars
Let # of 5-hr. courses be "42-x" ; income is 5*45(42-x)=9450-225x dollars
-------------
EQUATION:
income + income = 7020 dollars
135x + 9450-225x = 7020
-90x = -2430
x = 27 (# of 3-hr courses)
42-x = 15 (# of 5-hr courses)
==============
Cheers,
Stan H.